OpenStudy (anonymous):

Write an equation for the statement. Three more than twice a number x is fifteen.

OpenStudy (anonymous):

2x+3=15

OpenStudy (anonymous):

2x+3=15

OpenStudy (anonymous):

2x+3=15

OpenStudy (amistre64):

x = (15-3)/2 lol

OpenStudy (anonymous):

Wat? How do you get that? @amistre64

OpenStudy (anonymous):

Rearrange the equation for x?

OpenStudy (amistre64):

it is the voodoo the idoo soo well

OpenStudy (amistre64):

ever have a song stuck in your head that you cant quite recall the tune or the lyrics too? thats my life

OpenStudy (asnaseer):

best to work backwards, so final part says "is fifteen" so you know equation must be of the form\[... = 15\] "three more" means add 3 to whatever you have, giving\[... + 3 = 15\] "twice a number" - here you need to replace "a number" with a placeholder, lets call it x. So twice a number can be written as 2x, giving\[2x+3=15\]

OpenStudy (anonymous):

good answer

OpenStudy (anonymous):

Translating to Algebra you should read a little and translate a little. Three more than means that you're adding 3 to something so you have (something) +3 The question is what's 'something?' The something is twice a number (2x)+3 now comes the word that indicates you have an equation: the word 'is" which is the equal sign (2x)+3= finally we end with 15. (2x)+3=15 done. Note if it said 3 less than the translation would be (2x)-3=15 In both cases the parentheses can be removed because the is no number touching them on the left. Final answer: 2x+3 source: www.algebraINAflash.com